xzp - yzq= Kxyz by lagrange method


1
Expert's answer
2021-06-22T05:39:03-0400

This is Lagrange's equation


"Pp+Qq=R"




"P=xz, Q=-yz, R=Kxyz"

The auxiliary equation is,


"\dfrac{dx}{P}=\dfrac{dy}{Q}=\dfrac{dz}{R}"

"\dfrac{dx}{xz}=\dfrac{dy}{-yz}=\dfrac{dz}{Kxyz}"

"\dfrac{dx}{xz}=\dfrac{dy}{-yz}=>\dfrac{dx}{x}=-\dfrac{dy}{y}"

"\int \dfrac{dx}{x}=-\int\dfrac{dy}{y}"

"\ln x=-\ln y+\ln C_1"


"xy=C_1"

"\dfrac{dx}{xz}=\dfrac{dz}{Kxyz}"

"\dfrac{dx}{xz}=\dfrac{dz}{Kx(\dfrac{C_1}{x})z}"

"KC_1\dfrac{dx}{x}=dz"

"KC_1\ln x=z+C_2"

"Kxy\ln x-z=C_2"

Hence, the required general solution is given by


"\phi(xy, Kxy\ln x-z)=0"

where "\phi" is an arbitrary function.